Please show all work thank you
10). Adoor seen from above in the figures below) has hinges on the left hand side. The magkatudes of anforces arkaQual. Rank from smallest to greatest the torque applied by the force on each of the following situations. Explain how you determined your rank.


media%2F352%2F35217cea-022d-448a-9657-e1
Answer #1

10) Torque T = r*F*sin(theta)

r is the distance from axis of rotation to point of application of the force

F is the force acting

theta is the angle between r and F

A) r = 0
then torque TA = 0 N-m


B) theta lies between 0 and 90 deg

TB = r*F*sin(theta)

C) Torque is maximum since theta = 90 deg

TC= r*F

D) here theta lies between 0 and 180 deg

TD = r*F*sin(theta)

So the rank from smallest to greatest is

A < D < B < C

ADBC